f(x)=2x-3 and g(x)=x^2+2.
a) Find g(-4)
b) Show that gf(x)=4x^2-12x+11
c) Solve fg(x)=gf(x)

Answer by ikleyn(52797)

.
f(x)=2x-3 and g(x)=x^2+2.

a) Find g(-4).                      I sincerely hope you can substitute x= -4 into g(x)= x^2 + 2 and calculate g(-4).


b) Show that gf(x)=4x^2-12x+11.     g(f(x)) = (f(x))^2 + 2 = (2x-3)^2 + 2 = 4x^2 - 12x + 11.


c) Solve fg(x) = gf(x).            Hint:  Calculate f(g(x)) first. Then think.



----------------
Comment from student: Thanks i have solved the two questions!
im still stuck on the last one though, i have found fg(x) and gf(x), what would i do next?
----------------

My response:
If you just have the expressions for fg(x) and gf(x), you have to connect these two expressions
by the sign of equality "=" and to get an equation in this way.

Then you have to solve this equation for "x".

You are on the right track !